# Fusion-based Constitutive model (FuCe): Towards model-data augmentation in constitutive modelling

### v1

@@ Abstract
Constitutive modeling is crucial for engineering design and simulations to accurately describe material behavior. However, traditional phenomenological models often struggle to capture the complexities of real materials under varying stress conditions due to their fixed forms and limited parameters. While recent advances in deep learning have addressed some limitations of classical models, purely data-driven methods tend to require large datasets, lack interpretability, and struggle to generalize beyond their training data. To tackle these issues, we introduce "Fusion-based Constitutive model (FuCe): Towards model-data augmentation in constitutive modelling". This approach combines established phenomenological models with an ICNN architecture, designed to train on the limited and noisy force-displacement data typically available in practical applications. The hybrid model inherently adheres to necessary constitutive conditions. During inference, Monte Carlo dropout is employed to generate Bayesian predictions, providing mean values and confidence intervals that quantify uncertainty. We demonstrate the model’s effectiveness by learning two isotropic constitutive models and one anisotropic model with a single fiber direction, across six different stress states. The framework’s applicability is also showcased in finite element simulations across three geometries of varying complexities. Our results highlight the framework’s superior extrapolation capabilities, even when trained on limited and noisy data, delivering accurate and physically meaningful predictions across all numerical examples.
@@ 1 Introduction
In simulations of physical systems, constitutive modeling is essential for characterizing the mechanical behavior of materials. The accuracy of these simulations heavily depends on choosing the right constitutive model [1]. While some models are grounded in physics, most are primarily phenomenological. These models work well for small deformations in materials with linear behavior. However, when dealing with large deformations in materials like rubbers, polymers, foams, and biological tissues –which exhibit nonlinear and complex behaviors, developing accurate empirical or phenomenological models becomes significantly more difficult. Even with careful formulation and calibration, these models are limited by their fixed forms and the small number of parameters, restricting their ability to accurately represent a wide range of deformation levels and stress states. Machine learning based data-driven approach [2] including methods like Gaussian process regression [3, 4] or neural networks (NNs) [5, 6] presents a reasonable alternatives to overcome the time consuming task of formulating classical constitutive models and to improve the limited expressibility associated with these phenomenological models. The concept of using NNs in constitutive modeling was first presented in early 1990s [7]. The fundamental concept of machine-learning-based material modeling involves choosing a highly expressive model approach, characterized by a large number of learnable parameters, and then tuning these parameters using the available data. However, in this early phase, black-box approaches were mostly used, i.e., networks that do not take into account any physical principles and therefore only performs satisfactorily in the training regime. However, these models offer poorly generalization beyond the training data regime. Moreover, data-driven constitutive modeling requires a large number of strain-stress pairs. This presents two significant challenges, especially when relying on experimental data. First, it is impractical to explore the entire high-dimensional strain-stress space using standard mechanical tests, such as uni-/bi-axial tensile or bending tests. Second, complete tensorial constitutive models are difficult to develop because stress tensors cannot be directly measured; force measurements are only boundary-averaged projections of stress tensors. While multiscale simulations can generate extensive and detailed stress-strain data [8, 9], the computational cost for complex material systems is currently prohibitive. Therefore, it is crucial to learn material behavior directly from the force-displacement data that is practically available through mechanical testing.
Modern material model calibration primarily relies on full-field displacement data, such as that obtained via digital image correlation (DIC), combined with applied force data from load cells [10]. This calibration is typically performed using the finite element model updating (FEMU) [11] or virtual fields method (VFM) [12]. VFM directly solves for unknown material parameters by satisfying the weak form of momentum balance with measured displacement data, while FEMU iteratively optimizes model parameters until the simulated and measured displacement fields match. Recent efforts have demonstrated the estimation of stress fields using full-field displacement data in a model-free data-driven context [13, 14, 15]. In the model-based domain, physics-informed neural networks (PINNs) and their variants have shown promising results by first learning the forward solution to the mechanical boundary value problem and then estimating unknown parameters via gradient-based optimization [16, 17, 18, 19]. However, all these techniques, including FEMU, VFM, and PINNs, are limited to predefined constitutive models with a few unknown parameters. Such models are specific to particular materials and stress states and cannot be generalized beyond the calibration data.
@@
The limitations of the previously described frameworks can be mitigated by adopting a hybrid framework through data-model fusion [20]. These models integrate phenomenological and data-driven approaches. Various hybrid strategies exist for combining data-driven models with physics-based models to solve partial differential equations (PDEs). For instance, Chakraborty [21] introduced a transfer learning-based multi-fidelity physics-informed deep learning framework, which demonstrated efficiency but lacked interpretability. Our recent works [22, 23], also proposed an alternative framework that augments known physics with deep learning models for solving partial differential equations (PDEs) and stochastic differential equations (SDEs). These hybrid frameworks are found to be accurate, more interpretable and generalizes well to unseen environments.
In this research work, our objective is to develop a hybrid framework for constitutive modeling utilising existing phenomenological models along with the deep learning frameworks, using only the global reaction force and full-field displacement data. This approach is crucial due to the difficulty in availing stress strain data, practical problem of data scarcity, the difficulty in formulating accurate phenomenological models, and the critical need for generalization when a material’s behavior changes under varying external conditions. The salient features of the proposed FuCe approach are as follows:
@@
  • Model-data fusion: The proposed framework utilizes classical phenomenological constitutive models alongside deep learning models, which augment and correct the phenomenological models to allign with real material data.
  • Satisfies all necessary constitutive conditions: By employing an ICNN neural network with appropriate correction terms, the model satisfies all thermodynamic and physical feasibility conditions.
@@
  • Utilizes practically feasible force-displacement noisy data: The framework uses practically available full-field displacement and global reaction force data for training purposes, accounting for noise in the data.
  • Physics-informed loss: The training process is guided by a physics-informed loss function, compensating for the lack of direct stress/strain energy data.
@@
  • Approximate Bayesian inference: To quantify uncertainties in model predictions [24], a Bayesian approximation method called Monte Carlo Dropout [25] is used during the inference stage.
@@
The remainder of the paper is organized as follows. The background of constitutive models specifically hyperelasticity is formally described in Section 2. Detailed description on the problem statement is also presented at the end of Section 2. The proposed approach along with the algorithm is described in Section 3. Results and discussion of the numerical experiments that are carried out, are presented in Section 4. Finally, Section 5 presents the concluding remarks.
@@ 2.2 General criteria for hyper-elasticity
In elasticity, the goal of constitutive modeling is to establish a relationship between the strain at a material location and the corresponding stress. However, in hyperelastic constitutive modeling, instead of a direct mapping, the stress-strain relationship is derived through a potential $\psi\colon\mathcal{GL}^{+}(3)\to\mathbb{R}$, $\bm{F}\to\psi(\bm{F})$, which corresponds to the strain energy density stored in the body. This relationship can be expressed as:
@@
where $\mathbb{C}(\bm{F})$ denotes the (incremental) tangent modulus. Thus, the goal of hyperelastic constitutive modeling is to learn this strain energy potential $\psi(\bm{F})$. However, for any learned material model $\psi(\bm{F})$ to be mathematically plausible, it must also meet the following thermodynamic and physical requirements:
@@ 2.3 Classical Phenomenological models
    where $\lambda_{c}=\left(\frac{\tilde{I}_{1}}{3}\right)^{1/2}$, $\beta=\mathcal{L}^{-1}\left(\frac{\lambda_{c}}{\sqrt{N_{c}}}\right)$, and $\mathcal{L}^{-1}$ denotes the inverse Langevin function. The constants are set to $N_{c}=28$ (denoting the number of polymeric chain segments) and $c_{AB}\approx 3.7910$, the latter used to offset the energy density to zero at $\mathbf{F}=\mathbf{I}$ (since the Arruda–Boyce feature does not itself vanish for zero deformation).
@@
These models struggle to capture the complex behaviors of real materials due to their fixed form and restrictive functional dependence on invariants. The difficulty further escalates when trying to meet all the physical constraints outlined in Section 2.2. Additionally, each model has specific limitations under varying deformation levels and loading conditions. For example, the Neo-Hookean model is advantageous due to its simplicity and computational efficiency; however, it requires stringent conditions to achieve accurate results. The Mooney-Rivlin model performs well for small to medium strains, while the Gent model excels in predicting large strain loadings (up to 300%). The Isihara model is accurate for moderate stretch values but poorly estimates behavior under biaxial loading. The Ogden model is highly suitable for large deformation behavior; however, its drawback is the need for different sets of material parameters for each deformation mode.
@@ 2.4 Problem statement
Consider a 2D plate in a two-dimensional domain $\Omega$ with boundary $\partial\Omega$ is subjected to quasi-static loading. We apply Dirichlet boundary conditions on $\partial\Omega_{u}\subseteq\partial\Omega$ and Neumann boundary conditions on $\partial\Omega_{t}=\partial\Omega\setminus\partial\Omega_{u}$. The tests are conducted using displacement controlled setting, ensuring homogeneous Neumann boundary conditions and matching the forces applied by the testing equipment with the reaction forces at the fixed boundary. The available data includes displacement measurements $\{\bm{u}^{t}_{r}:r=1,...,n_{d};t=1,...,n_{l}\}$ at $n_{d}$ points and net reaction force measurements $\{\bm{R}^{t}_{\beta}:\beta=1,...,n_{\beta};t=1,...,n_{l}\}$ on some boundary segments, collected over $n_{l}$ load steps. The net reaction forces on segments of $\partial\Omega_{u}$ (e.g., each loaded side of the specimen) are assumed to be known, simulating measurements from load cells. With this setup, the overall objective is to discover the underlying constitutive law of the unknown material. We have prior knowledge that the material exhibits hyperelastic behavior, and a potential constitutive model, $\psi_{\text{known}}(\bm{F})$ is also available. However, as is often the case, $\psi_{\text{known}}(\bm{F})$ is approximate, and the true constitutive relation takes the following form
$\psi(\bm{F})=\psi_{\text{known}}(\bm{F})+\psi_{\text{missing}}(\bm{F}).$ (21)
cannot rely solely on phenomenological models. Since the constitutive model $\psi_{\text{known}}(F)$ is not accurate, the obtained stress field will also be inaccurate, leading to an erroneous force field solution. With this setup, the objective is to learn the missing physics, $\psi_{\text{missing}}(\bm{F})$ by using the force-displacement data as described before.
@@ 3.2 ICNN Architecture
Having discussed the fundamental constraints and a systematic procedure for satisfying the same, we proceed to discussing the architecture of the Input Convex Neural Network (ICNN), which is a key component of the proposed approach. The inputs to the ICNN are modified versions of various invariants of $\bm{C}$, as detailed in Section.2.3. The architecture consists of repeated convex layers, each with 64 neurons in the hidden layer. It incorporates skip connections from the input to the hidden layer after each convex layer. This sequence is repeated three times. Convexity is ensured by maintaining positive weights, achieved by applying the softplus function to the weight matrix prior to its use, as shown below:
@@
respectively, for all $z\in\mathbb{R}$. Here, $\mathcal{\gamma}_{1}>0$ and $\mathcal{\gamma}_{2}>0$ are hyperparameters. The softplus function guarantees $C^{\infty}$ continuity of the strain energy density and its derivatives, ensuring that stress and tangent are sufficiently smooth and continuous. The function $\sigma_{1}$ is taken as the squared softplus to address the issue of the vanishing second derivative. This squaring not only facilitates the calculation of the tangent of the strain energy density but also prevents vanishing gradients during backpropagation in neural network training. Additionally, skip connections are incorporated to mitigate the issues of vanishing gradients, overfitting, and accuracy degradation (or saturation) as the neural network architecture deepens. At the output of the neural network we get the missing component of the strain energy density.
@@ 3.3 Training
As discussed in Section 2.4, we have nodal displacement data represented as $u^{t}(r,j)$, where $t=\{1,2,\ldots,n_{t}\}$ denotes the load step, $r=\{1,2,\ldots,n_{d}\}$ indexes the nodal points, and $j=\{1,2\}$ indicates the direction of displacement. We define the degree of freedom at the $r^{\text{th}}$ node in the $j^{\text{th}}$ direction as $\mathcal{D}_{f}=(r,j)$. This encompasses both $\mathcal{D}_{f}^{\text{free}}$: degrees of freedom within the boundary where displacement is not fixed, and $\mathcal{D}_{f}^{\text{fixed}}$: degrees of freedom at the boundaries where displacement is constrained. Additionally, we have reaction force data $R_{\beta}$ available at selected boundary nodes $\beta=1,2,\ldots,n_{\beta}$.
@@
Figure 1: Training Setup: A square plate with a hole in the bottom-left corner is subjected to asymmetric biaxial tension under displacement control. The noisy data collected from full-field displacements and reaction forces is used to train the proposed constitutive model.
@@
From the displacement field $\bm{u}(\bm{X})$, we obtain the deformation gradient following Eq. (31), modified invariants of which are input into the ICNN to yield the strain energy density potential. Using Eq. (11), we derive the stress field. In the absence of stress field data, we use physics-informed loss in form of the linear momentum conservation equation to inform and guide the training process. Note that although we use finite element method in this work, other techniques including finite volume and finite difference can also be used.
@@
Although we have used finite element implemented using differentiable finite element simulator, the overall approach can be seamlessly integrated with other numerical techniques such as finite volume method and finite difference method.
To optimize the learnable parameters, we derive the loss function using Eq. (35):
@@
In this optimization problem, although the network is convex with respect to its input, it is non-convex concerning the trainable parameters. Additionally, direct stress or strain energy labels are absent, so we use global reaction force data for training. These factors lead to an optimization problem with multiple solutions (multiple local minima), making the solution highly sensitive to initialization. To address this challenge, we trained multiple ICNNs independently with different random initializations. The model with the minimum loss value among all the trained models is selected and used for prediction. The training algorithm is shown in Algorithm 1.
Input: Point-wise displacement data $U=\{\bm{u}^{t}_{r,j}\},e\in\mathbb{R}^{2},r=1,\ldots,n_{d},t=1,\ldots,n_{l}$
Input: Global reaction forces $R^{\beta}=\{R^{t}_{r,j},\beta=1,\ldots,n_{i},t=1,\ldots,n_{t}\}$
1 Randomly initialize ICNN parameters $\theta$ and the Adam optimizer with a learning rate scheduler;
2 for $e=1,\ldots,epochs$ do
3       $\mathcal{L}\leftarrow 0$ ;
4        // Initialize loss for current epoch

### v2

@@ Abstract
Constitutive modelling is crucial for engineering design and simulations to accurately describe material behavior. However, traditional phenomenological models often struggle to capture the complexities of real materials under varying stress conditions due to their fixed forms and limited parameters. While recent advances in deep learning have addressed some limitations of classical models, purely data-driven methods tend to require large datasets, lack interpretability, and struggle to generalize beyond their training data. To tackle these issues, we introduce "Fusion-based Constitutive model (FuCe): Towards model-data augmentation in constitutive modelling". This approach combines established phenomenological models with an ICNN architecture, designed to train on the limited and noisy force-displacement data typically available in practical applications. The hybrid model inherently adheres to necessary constitutive conditions. During inference, Monte Carlo dropout is employed to generate Bayesian predictions, providing mean values and confidence intervals that quantify uncertainty. We demonstrate the model’s effectiveness by learning two isotropic constitutive models and one anisotropic model with a single fibre direction, across six different stress states. The framework’s applicability is also showcased in finite element simulations across three geometries of varying complexities. Our results highlight the framework’s superior extrapolation capabilities, even when trained on limited and noisy data, delivering accurate and physically meaningful predictions across all numerical examples.
@@ 1 Introduction
In simulations of physical systems, constitutive modelling is essential for characterizing the mechanical behaviour of materials. The accuracy of these simulations heavily depends on choosing the right constitutive model [1]. While some models are grounded in physics, most are primarily phenomenological. These models work well for small deformations in materials with a linear behaviour. However, when dealing with large deformations in materials like rubbers, polymers, foams, and biological tissues – which exhibit non-linear and complex behaviours, developing accurate empirical or phenomenological models becomes significantly more difficult. Even with careful formulation and calibration, these models are limited by their fixed forms and the small number of parameters, restricting their ability to accurately represent a wide range of deformation levels and stress states. Machine learning-based data-driven approach [2] including methods like Gaussian process regression [3, 4] or neural networks (NNs) [5, 6] presents reasonable alternatives to overcome the time-consuming task of formulating classical constitutive models and to improve the limited expressibility associated with these phenomenological models. The concept of using NNs in constitutive modelling was first presented in the early 1990s [7]. The fundamental concept of machine-learning-based material modeling involves choosing a highly expressive model approach, characterized by a large number of learnable parameters, and then tuning these parameters using the available data. However, in this early phase, black-box approaches were mostly used, i.e., networks that do not take into account any physical principles and therefore only perform satisfactorily in the training regime. However, these models offer poor generalization beyond the training data regime. In addition, a large dataset of stress-strain pairs is required to train the constitutive models. This presents two new issues when using experimental data. First, it is impractical to investigate the complete high-dimensional stress-strain space via conventional mechanical experiments, such as uniaxial or biaxial tensile and bending tests. Second, constructing comprehensive constitutive tensorial models is challenging since stress tensors are not directly measurable, and force measurements only represent boundary average projections of stress tensors, as highlighted in [8]. Although multiscale simulations can produce large, detailed datasets of stress-strain relationships [9, 10], their computational demands for complex material systems remain prohibitive. Therefore, it is crucial to learn material behaviour directly from the force-displacement data that is practically available through mechanical testing.
There have been attempts towards discovering constitutive relations from force-displacement data. For example, EUCLID and its variants [11, 8, 12] have recently been proposed that eliminate the requirement of stress data in model training. Instead, it leverages displacement and force data to learn the material’s strain energy density function, enabling the formulation of hyperelastic constitutive laws directly from experimental observations. This method is especially beneficial in situations where stress measurements are unfeasible or susceptible to inaccuracies. Of particular interest is the NN-Euclid [11] framework that utilizes the input constraint neural network (ICNN) to learn the constitutive relation from force-displacement data. However, NN-Euclid [11] does not exploit known (but approximate) information about the material model, that often is a priori available. We hypothesize that exploiting such prior information can help the neural network to learn from sparse data and extrapolate better, and model-data fusion is potentially the way forward in constitutive modelling.
@@
In model-data fusion, the high-level idea is to integrate phenomenological and data-driven models [13]. Along similar line, hybrid strategies already exist for combining data-driven models with physics-based models to solve partial differential equations (PDEs). For instance, Chakraborty [14] introduced a transfer learning-based multi-fidelity physics-informed deep learning framework, which demonstrated efficiency but lacked interpretability. Our recent works [15, 16], also proposed an alternative framework that augments known physics with deep learning models for solving PDEs and stochastic differential equations (SDEs). These hybrid frameworks are found to be accurate, more interpretable and generalize well to unseen environments. However, contrary to these work, the goal in this paper is not to solve PDEs/SDEs using data-physics fusion; instead, the objective is to develop a hybrid framework for constitutive modelling utilising existing phenomenological models along with the deep learning frameworks. Additionally, similar to [11], we want to train the model using only the global reaction force and full-field displacement data. This is crucial due to the difficulty in availing stress-strain data, the practical problem of data scarcity, the difficulty in formulating accurate phenomenological models, and the critical need for generalization when a material’s behaviour changes under varying external conditions.
In this research work, we propose FuCe, a novel approach rooted in model-data augmentation for learning constitutive learning from global force displacement data. FuCe extends the NN-Euclid proposed in [11] by incorporating a known (but approximate) model into the overall framework. Therefore, all the advantages inherent in NN-Euclid also hold for FuCe while allowing learning from sparse data and better generalization beyond the training regime. Additionally, we also incorporate Monte Carlo-dropout (MC-dropout) into the architecture, making it uncertainty-aware. Overall, the salient features of the proposed FuCe approach are as follows:
@@
  • Model-data fusion: The proposed framework utilizes classical phenomenological constitutive models alongside deep learning models, which augment and correct the phenomenological models to align with real material data.
  • Satisfies all necessary constitutive conditions: Similar to NN-Euclid [11], FuCe satisfies all thermodynamic and physical feasibility conditions.
@@
  • Utilizes practically feasible force-displacement noisy data: As the proposed FuCe builds on top of NN-Euclid, it allows learning from practically available full-field displacement and global reaction force data for training purposes, accounting for noise in the data.
@@
  • Approximate Bayesian inference: To quantify uncertainties in model predictions [17], a Bayesian approximation method called Monte Carlo Dropout [18] is used during the inference stage. This makes the model uncertainty-aware and the uncertainty estimated can be used for decision-making, if needed.
@@
The remainder of the paper is organized as follows. The background of constitutive models specifically hyperelasticity is formally described in Section 2. A detailed description of the problem statement is also presented at the end of Section 2. The proposed approach along with the algorithm is described in Section 3. Results and discussion of the numerical experiments that are carried out are presented in Section 4. Finally, Section 5 presents the concluding remarks.
@@ 2.2 General criteria for hyper-elasticity
In elasticity, the goal of constitutive modelling is to establish a relationship between the strain at a material location and the corresponding stress. However, in hyperelastic constitutive modelling, instead of a direct mapping, the stress-strain relationship is derived through a potential $\psi\colon\mathcal{GL}^{+}(3)\to\mathbb{R}$, $\bm{F}\to\psi(\bm{F})$, which corresponds to the strain energy density stored in the body. This relationship can be expressed as:
@@
where $\mathbb{C}(\bm{F})$ denotes the (incremental) tangent modulus. Thus, the goal of hyperelastic constitutive modelling is to learn this strain energy potential $\psi(\bm{F})$. However, for any learned material model $\psi(\bm{F})$ to be mathematically plausible, it must also meet the following thermodynamic and physical requirements:
@@ 2.3 Classical Phenomenological models
    where $\lambda_{c}=\left(\frac{\tilde{I}_{1}}{3}\right)^{1/2}$, $\beta=\mathcal{L}^{-1}\left(\frac{\lambda_{c}}{\sqrt{N_{c}}}\right)$, and $\mathcal{L}^{-1}$ denotes the inverse Langevin function. Following [11], the constants are set to $N_{c}=28$ (denoting the number of polymeric chain segments) and $c_{AB}\approx 3.7910$, the latter used to offset the energy density to zero at $\mathbf{F}=\mathbf{I}$ (since the Arruda–Boyce feature does not itself vanish for zero deformation).
@@
These models struggle to capture the complex behaviours of real materials due to their fixed form and restrictive functional dependence on invariants. The difficulty further escalates when trying to meet all the physical constraints outlined in Section 2.2. Additionally, each model has specific limitations under varying deformation levels and loading conditions. For example, the Neo-Hookean model is advantageous due to its simplicity and computational efficiency; however, it requires stringent conditions to achieve accurate results. The Mooney-Rivlin model performs well for small to medium strains, while the Gent model excels in predicting large strain loadings (up to 300%). The Isihara model is accurate for moderate stretch values but poorly estimates behaviour under biaxial loading. The Ogden model is highly suitable for large deformation behaviour; however, its drawback is the need for different sets of material parameters for each deformation mode.
@@ 2.4 Problem statement
Material characterisation involves understanding the constitutive behaviour of the material under various loading conditions. In practice, reaction forces and displacement measurements are often used as primary data, however, these measurements are typically sparse and are susceptible to inaccuracies. This poses challenges in accurately discovering material behaviour. On the other hand, phenomenological models are often approximate and fail to capture the material’s true response accurately. We hypothesize that these challenges can be addressed by augmenting known phenomenological models with machine learning-based models.
To formally define the problem, consider a material domain $\Omega$ undergoing quasi-static loading in displacement-controlled testing. The configuration incorporates Dirichlet and Neumann boundary conditions, with the applied forces corresponding to reaction forces at constrained boundaries. The data comprises limited displacement measurements, and net reaction forces collected over load steps on specific boundary segments, emulating load cell measurements. With this setup, the objective is to discover the underlying constitutive law of the unknown material. Note that [11] have already tackled this problem setup; however, the key difference here resides in the fact that we have prior (approximate) knowledge that the material exhibits hyperelastic behaviour, and a potential constitutive model is also available. Since the constitutive model is approximate, the obtained stress field will also be inaccurate, leading to an erroneous force field solution. With this setup, the objective is to learn a hybrid model that fuses the known phenomenological model with the data-driven neural network.
@@ 3.2 ICNN Architecture
Having discussed the fundamental constraints and a systematic procedure for satisfying the same, we proceed to discussing the architecture of the Input Convex Neural Network (ICNN), which is a key component of the proposed approach. Similar to previous work [11], the inputs to the ICNN are modified versions of various invariants of $\bm{C}$, as detailed in Section.2.3. The architecture consists of repeated convex layers, each with 64 neurons in the hidden layer. It incorporates skip connections from the input to the hidden layer after each convex layer. This sequence is repeated three times. Convexity is ensured by maintaining positive weights, achieved by applying the softplus function to the weight matrix prior to its use, as shown below:
@@
respectively, for all $z\in\mathbb{R}$. Here, $\mathcal{\gamma}_{1}>0$ and $\mathcal{\gamma}_{2}>0$ are hyperparameters. The softplus function ensures $C^{\infty}$ continuity of the strain energy density and its derivatives, providing smooth and continuous stress and tangent operators. We consider the function $\sigma_{1}$ as the squared softplus to handle the vanishing second derivative problem. This squaring not only helps to compute the tangent of the strain energy density but also stops disappearing gradients during backpropagation in neural network training (see [32]). Additionally, skip connections are incorporated to mitigate the issues of vanishing gradients, overfitting, and accuracy degradation (or saturation) as the neural network architecture deepens, see reference [33]. At the output of the neural network we get the missing component of the strain energy density.
@@ 3.3 Training
Consider we have nodal displacement data represented as $u^{t}(r,j)$, where $t=\{1,2,\ldots,n_{t}\}$ denotes the load step, $r=\{1,2,\ldots,n_{d}\}$ indexes the nodal points, and $j=\{1,2\}$ indicates the direction of displacement. We define the degree of freedom at the $r^{\text{th}}$ node in the $j^{\text{th}}$ direction as $\mathcal{D}_{f}=(r,j)$. This encompasses both $\mathcal{D}_{f}^{\text{free}}$: degrees of freedom within the boundary where displacement is not fixed, and $\mathcal{D}_{f}^{\text{fixed}}$: degrees of freedom at the boundaries where displacement is constrained. Additionally, we have reaction force data $R_{\beta}$ available at selected boundary nodes $\beta=1,2,\ldots,n_{\beta}$.
@@
Figure 1: Training Setup: A square plate with a hole in the bottom-left corner is subjected to asymmetric biaxial tension under displacement control. The noisy data collected from full-field displacements and reaction forces is used to train the proposed constitutive model. This setup is the same as [11] so as to illustrate the advantage of model-physics fusion.
@@
From the displacement field $\bm{u}(\bm{X})$, we obtain the deformation gradient following Eq. (30), modified invariants of which are input into the ICNN to yield the strain energy density potential. Using Eq. (11), we derive the stress field. In the absence of stress field data, we use physics-informed loss in the form of the linear momentum conservation equation to inform and guide the training process. Note that although we use the finite element method in this work, other techniques including finite volume and finite difference can also be used.
@@
Although we have used finite element implemented using a differentiable finite element simulator, the overall approach can be seamlessly integrated with other numerical techniques such as finite volume method and finite difference method. For more detials about the Eq. (34) and Eq. (33) interested readers can refer to [11].
To optimize the learnable parameters, we employ a loss function similar to the one used in [11], which can be expressed as:
@@
Note that unlike [11], our overall model also includes the known phenomenological model. Note that although the network (corresponding to the missing physics) exhibits convexity with respect to its input but is non-convex in relation to the trainable parameters. Moreover, due to the absence of direct stress or strain energy labels, we utilize global reaction force data for training. These factors lead to an optimization problem with multiple solutions (multiple local minima), making the solution highly sensitive to initialization. To address this challenge, we trained multiple networks independently with different random initialisations. The model with the lowest loss value among all the trained candidates is selected and subjected to Monte Carlo dropout during prediction to account for uncertainty. The details of the training process are described in Algorithm 1.
1
Input: Displacement data $U=\{\bm{u}^{t}_{r,j}\},$ where $r=1,\ldots,n_{d},$ $j=1,2,$ $t=1,\ldots,n_{l}$; and global reaction forces $R^{\beta}=\{R^{t}_{r,j},\beta=1,\ldots,n_{i},t=1,\ldots,n_{t}\}$
2
3
4Initialize ICNN parameters $\bm{\theta}$ and Adam optimizer;
5
6for $e=1,\ldots,epochs$ do
7       Set $\mathcal{L}\leftarrow 0$ ;
8        // Initialize epoch loss
9